Range of projectile motion =500mtime=10secondfind maximum height?

Grade:12th pass

1 Answers

Vikas TU
7 years ago
500 = u^2*sin(2theta)/g................(1)
and
10 = u^2*(sin(theta))^2/2g..................(2)
solving both eqns. we get,
ucos(theta) = 50 
and
usin(theta) = 50
we  get u = root(5000)
Maximum height = u^2/2g = > 5000/20 = > 250 meter.